# Tractable Probabilistic Models for Investment Planning

### v1

@@ 3.2 TPM-Based Pipeline for Probabilistic Energy System Planning
To demonstrate that TPMs could solve the Capacity/Generation/Storage Expansion problems, we propose a pipeline organized into three layers as shown in Figure 2. The first layer is a realistic simulator representing a generation system design scenario, in which the decisions (based on hardware specifications), in conjunction with prototypical household load shapes and stochastic environmental variables, will determine whether the system will fulfill its demands successfully or suffer shortfalls. The simulator is a scenario generator that produces large numbers of daily realizations. The second layer performs probabilistic learning using SPNs, which capture the conditional probability of shortfall events for each possible decision. The third layer couples these probabilistic outputs with an MILP engine, enabling risk-aware design by imposing chance constraints on shortfall probability.
@@ 3.2.2 Layer 2: SPN Training and Inference
The second layer learns a tractable probabilistic representation of the relationship between design choices, environmental factors, and adequacy constraints using the labeled dataset generated in Layer 1. We train an SPN to approximate the joint density $p(X,\Xi,Y)$, where $X$ represents design and control variables, $\Xi$ the stochastic environmental inputs, and $Y$ the adequacy label (shortfall or no-shortfall). Once trained, the SPN supports exact density queries such as the conditional probability of system adequacy given a specific design, $\mathbb{P}(Y=1\mid X=x)$, and marginal probabilities across subsets of variables. Because SPNs permit inference in polynomial time in the network size, the resulting model enables the evaluation of chance constraints over arbitrarily large scenario spaces without explicit enumeration. This layer bridges the data-driven simulation environment with the optimization stage by providing an interpretable and tractable probabilistic surrogate for system reliability.
@@ 3.2.3 Layer 3: MILP Integration and Optimization
The third layer integrates the trained SPN into a mixed-integer linear programming (MILP) framework to perform risk-aware capacity, generation, and storage planning. Notice that in the chance constraint (5), one could consider the probability of “inadequacy”, i.e., the event where the demand exceeds generation by more than some tolerance $\varepsilon$. The positive part of demand with the generation subtracted is sometimes also known as the (instantaneous) shortfall. The optimization seeks a configuration $x$ minimizing total investment and operational cost while ensuring that the probability of inadequacy remains below a predefined threshold, throughout the planning horizon. For practical purposes, in our methodology, we will reduce the objective to only the investment component:
@@ 4.1 Probability Landscape Comparison
To evaluate the SPN-based methods’ ability to learn the true probability distribution over system configurations, we compare their probability landscapes against the empirical estimation of our probability of interest (20) across different levels of data availability. We aimed to check whether high‑probability regions, ridges, and exclusion zones are recovered in the right places, how estimates evolve with more data (convergence), and where systematic biases or overconfidence may persist.
Figure 3: Comparison between empirical and SPN-estimated shortfall ratios (SPN and SPN-Max) for the dataset percentage. The left column shows the best empirical estimation we could have with the available data. The middle column shows the fitted SPN estimation, and the right column shows the SPN-Max, which introduces error with the sum node approximations. As data accessibility grows, the SPN increasingly recovers the diagonal pattern that shows a transition boundary between high and low probability of inadequacy regimes.
@@
The key analytical value of the grid layout in Figure 3 lies in enabling two critical comparisons simultaneously: horizontal comparison within each row reveals whether the SPN-based methods accurately capture the probability patterns of the ground truth for a given data availability scenario, while vertical comparison across rows shows how these probability estimates evolve and potentially converge toward ground truth as more training data becomes available. Ideally, the three heatmaps in each row should exhibit similar color patterns, with bright yellow regions (high probability) appearing in the same locations across all three methods, indicating that the SPN has correctly learned which configurations correspond to a low likelihood of a shortfall event.
@@
The results show that SPN and SPN-Max estimates progressively align more closely with the empirical reference patterns, reflecting the improvement in model accuracy with increased training data. The comparison between SPN (expectation over all paths) and SPN-Max (maximum probability path) also reveals important characteristics about the network structure: significant differences between these two suggest that the probability mass is distributed across multiple paths, while near-identical patterns indicate the network relies heavily on a dominant path, demonstrating that the approximation in SPN-Max comes at the cost of imprecision in cases where there is not a dominant path in sum nodes. This experiment also allows us to check whether poor optimization decisions stem from inaccurate probability modeling (misaligned heatmap patterns) or from issues in the optimization procedure itself (correct probability patterns but suboptimal configuration selection).
@@
Figure 4: Configuration space exploration across data sparsity levels (30 iterations per percentage). Point clouds show each method’s configuration choices with capacity-based axes (primary) and unit counts (secondary). A gold star indicates the ground truth optimum. Red dashed boundary marks the adequacy constraint ($\mathbb{P}(\text{shortfall})=\varepsilon$, where $\varepsilon=5\%$) based on ground truth. Configurations outside this boundary fail the adequacy requirement. Note that these methods may appear safe by their own estimates, yet select configurations beyond the true safety boundary, revealing calibration overconfidence.
@@ 4.2 Reliability of SPN-Based Investment Planning under Data Availability Constraints
    s.t. $G(x,\xi)=0,\quad\text{with}~\xi\sim\hat{\Xi}$
    We assumed that optimizing a larger set is prohibitive, and the objective aims to minimize the investment in a configuration that satisfies adequacy. This formulation tends to underestimate risk because of its limited stochastic coverage.
@@
  • SPN: Uses the fitted SPN (using the available data sampled for the specific iteration) to directly estimate the probability of adequacy violations. This estimation cannot be the input of an optimizer, but it is still a point of reference for what the best solution the SPN-Max approximation could yield.
  • SPN-Max: This value follows the stimation proposed in [nemecek2024generating] to have an MILP-compatible function. This approach enables embedding SPNs within a tractable optimization pipeline, allowing explicit enforcement of chance constraints during decision-making.
@@
This experiment reveals some important insights that validate our method but also show some important considerations. First, it illustrates the progressive improvement of the SPN’s reliability with larger dataset sizes. A limited number of samples allows fitting an SPN roughly; it will learn some high-level patterns that might allow finding a feasible region to make a decision, but as it receives more data, it will undergo a process of refinement that could damage the quality of the estimation (like in the case of $50\%$). The SPN-Max approximation depends directly on how well the SPN is fitted, which will be posteriorly encoded in the MILP. However, our test showed that the error introduced by the approximation of the sum node might have a regularization effect (causing the decision to be conservative), which, as a consequence, keeps the decision consistently within the allowance constraint throughout the different data availability scenarios, which, in our experiment, is essential.
@@
Comparing against the GEP10 and GEP50 approaches, there are some advantages of using SPNs. In our experimental setup, the traditional GEP approaches depend a lot on the quality of the selected scenarios, which we test by selecting them from the available samples on each iteration of our experiment. When choosing a few scenarios, there is more uncertainty on fulfilling the adequacy constraint, which we evidence in Figure 4 as GEP10 shows large variability among the iterations, and with no proper guarantees of preventing shortfall. Switching to GEP50, considering a larger set of reference scenarios improves the consistency of the decision; however, it doesn’t offer certainty about respecting the adequacy either. In both cases, it becomes a matter of the quality of the available scenarios.
@@
We noticed that GEP10 shows high variability in the cost of the selected solution, which is just a consequence of its dependence on the scenario selection and the limitation on the number of scenarios one can consider ($K$). The case of GEP50 shows an expected outcome: solutions would tend to be more expensive than the optimal one, but they show the lowest deviation from the optimal cost. However, the average solution lies on the boundary of adequacy, meaning that any error in the selection or simulation could move the solution outside of the safe area. Finally, the SPN methods tend to be conservative in the decision, a product of the quality of the fitting process and the error introduced by the approximation (in the case of SPN-Max). Although more expensive, SPN methods show more consistency throughout data availability scenarios and random samples. This is a clear advantage, considering that the operation cost component $C^{\mathrm{op}}(x,\xi)$ in (4) would be negatively affected by shortfall events. Thus, we argue that although more expensive, TPM-based methods to solve the chance-constrained Capacity/Generation/Storage expansion problem offer a safer and cheaper (in the long term) alternative to traditional methods.
@@ 4.3 Training Details
We trained our SPNs using the SPFlow library [Molina2019SPFlow], which supports automatic structure learning and exact inference for tractable probabilistic models through queries. The structure learning procedure relies on hierarchical clustering and recursive variable partitioning to identify conditional independencies within the data. We optimized the hyperparameters using WandB [wandb] through Bayesian search over 500 trials, per percentage of the total dataset, with model selection based on ground-truth log-likelihood on the validation set. We standardized and discretized features depending on the leaf node type. Individual runs required from a few seconds to several minutes, depending on the dataset size and model complexity. Early stopping based on validation likelihood was employed to prevent overfitting. To ensure full reproducibility, we plan to make all code, configuration files, and instructions available to replicate the experiments.
@@ 5 Conclusion
This work presents a probabilistic framework for investment planning in power systems under uncertainty, based on Sum–Product Networks (SPNs), by embedding these SPNs within a mixed-integer linear programming (MILP) formulation. Based on existing work on Tractable Probabilistic Models (TPMs), we showed that we can incorporate probabilistic reasoning into chance-constrained optimization problems without explicit scenario enumeration, redefining the traditional capacity/generation/storage scenario expansion problem by integrating a probabilistic adequacy constraint.
Our experimental results on a representative generation–storage planning case study indicated that SPN-based models can accurately capture adequacy probabilities and maintain reliable performance even under limited data availability. Compared to traditional scenario-based methods, the SPN formulation yielded more consistent and conservative investment decisions, mitigating shortfall risk at moderate computational cost. These findings highlight the potential of tractable probabilistic models as a foundation for scalable, data-driven, and reliability-aware expansion planning in future low-carbon power systems, which rely on stochastic non-dispatchable sources.
This work demonstrates that tractable probabilistic models provide a promising and interpretable path toward integrating uncertainty quantification and optimization in long-term power system planning.
@@ 6 Limitations
While the proposed framework exhibits strong potential, there are some significant limitations to highlight that could enable further research. First, the SPN-Max linearization introduces an approximation that may lead to conservative or locally imprecise density estimation, yielding a calibration problem and potentially introducing overconfidence when multiple probabilistic paths contribute significantly to the overall likelihood. Future work will explore alternative formulations, other probabilistic circuits, or hybrid approaches to reduce this approximation error. Additionally, the quality of probabilistic inference depends on the representativeness of the training data. Real-world datasets may exhibit bias or insufficient coverage of rare events, which could affect reliability estimation. Incorporating active learning or uncertainty-aware sampling could improve robustness in such settings. Although the TPM formulation scales more favorably than scenario-based methods, embedding large SPNs within MILP solvers can pose computational challenges for full-scale systems if the learned structure grows larger.

### v2

@@ 3.2 TPM-Based Pipeline for Probabilistic Energy System Planning
To demonstrate that TPMs can solve the Capacity/Generation/Storage Expansion problems, we propose a pipeline organized into three layers as shown in Figure 2. The first layer is a realistic simulator representing a generation system design scenario, in which decisions (based on hardware specifications) are made in conjunction with prototypical household load shapes and stochastic environmental variables, determining whether the system will fulfill its demands successfully or suffer shortfalls. The simulator is a scenario generator that produces large numbers of daily realizations. The second layer performs probabilistic learning using SPNs, which capture the conditional probability of shortfall events for each possible decision. The third layer couples these probabilistic outputs with an MILP engine, enabling risk-aware design by imposing chance constraints on shortfall probability.
@@ 3.2.2 Layer 2: SPN Training and Inference
The second layer learns a tractable probabilistic representation of the relationship between design choices, environmental factors, and adequacy constraints using the labeled dataset generated in Layer 1. We train an SPN to approximate the joint density $p(X,\Xi,Y)$, where $X$ represents design and control variables, $\Xi$ the stochastic environmental inputs, and $Y$ the adequacy label (shortfall or no-shortfall). Once trained, the SPN supports exact density queries such as the conditional probability of system adequacy given a specific design, $\mathbb{P}(Y=1\mid X=x)$, and marginal probabilities across subsets of variables. Because SPNs permit inference in polynomial time with respect to the network size, the resulting model enables the evaluation of chance constraints over arbitrarily large scenario spaces without explicit enumeration. This layer bridges the data-driven simulation environment with the optimization stage by providing an interpretable and tractable probabilistic surrogate for system reliability.
@@ 3.2.3 Layer 3: MILP Integration and Optimization
The third layer integrates the trained SPN into an MILP framework to perform risk-aware capacity, generation, and storage planning. Notice that in the chance constraint (5), one could consider the probability of “inadequacy”, i.e., the event where the demand exceeds generation by more than some tolerance $\varepsilon$. The positive part of demand with the generation subtracted is sometimes also known as the instantaneous shortfall. The optimization seeks a configuration $x$ that minimizes total investment and operational costs while ensuring that the probability of inadequacy remains below a predefined threshold throughout the planning horizon. For practical purposes, in our methodology, we will reduce the objective to only the investment component:
@@ 4.1 Probability Landscape Comparison
To evaluate the ability of SPN-based methods to learn the true probability distribution over system configurations, we compare their probability landscapes with the empirical estimation of our probability of interest (20) across different levels of data availability. We aimed to verify whether high-probability regions, ridges, and exclusion zones are accurately recovered in the correct locations, how estimates evolve with additional data (convergence), and where systematic biases or overconfidence may persist.
Figure 3: Comparison between empirical and SPN-estimated shortfall ratios (SPN, SPN-Max, and SPN-Piecewise): In order, the Empirical column shows the best empirical estimation possible with respect to access to more samples. The SPN column displays the learned approximation. The SPN-Max column shows the first proposed approximation (SPN-Max). The SPN-Piecewise column displays the improved version of our proposed approximation. We highlight that our best method is capable of capturing the pattern of the best empirical approximation. Although SPN-Max captures the pattern as well, the error scales with the size of the SPN, something that is minimized with SPN-Piecewise while remaining computable using MILP.
@@
The key analytical value of the grid layout in Figure 3 lies in enabling two critical comparisons simultaneously: horizontal comparison within each row reveals whether the SPN-based methods accurately capture the probability patterns of the best empirical estimator for a given data availability scenario, while vertical comparison across rows shows how these probability estimates evolve and potentially converge toward ground truth as more training data becomes available. Ideally, the heatmaps in each row should exhibit similar color patterns, with bright yellow regions (indicating high probability) appearing in the same locations across all methods, suggesting that the SPN has correctly learned which configurations correspond to a low likelihood of a shortfall event.
@@
The results show that SPN progressively aligns more closely with the empirical reference patterns, reflecting the improvement in the model’s ability to represent the desired density as the training data increases. The comparison between SPN (expectation over all paths) and SPN-Max (maximum probability path) reveals important characteristics about the network structure: significant differences between these two suggest that the probability mass is distributed across multiple paths, while near-identical patterns indicate the network relies heavily on a dominant path, demonstrating that the approximation in SPN-Max comes at the cost of imprecision in cases where there is not a dominant path in sum nodes. As a consequence, the error grows with the number of nodes of the SPN. The SPN-Piecewise approximation significantly reduces the error. We can conclude that poor optimization decisions stem from inaccurate probability modeling (misaligned heatmap patterns) or from issues in the optimization procedure itself (correct probability patterns but suboptimal configuration selection).
@@
Figure 4: Configuration space exploration across data sparsity levels (30 iterations per percentage). Point clouds show each method’s configuration choices with capacity-based axes (primary) and unit counts (secondary). A gold star indicates the ground truth optimum. Red dashed boundary marks the region where the adequacy constraint ($\mathbb{P}(\text{shortfall})=\varepsilon$, where $\varepsilon=5\%$) is respected based on ground truth. Configurations outside this boundary fail the adequacy requirement. Note that these methods may appear safe based on their own estimates, yet select configurations that exceed the true safety boundary, revealing calibration overconfidence. We added some jitter to the points to show the clusters of runs, but they concentrate around the same solution.
@@ 4.2 Reliability of SPN-Based Investment Planning under Data Availability Constraints
    s.t. $G(x,\xi)\geq 0,\quad\text{with}~\xi\sim\hat{\Xi}$
    We assumed that optimizing a larger set is prohibitive, and the objective aims to minimize the investment in a configuration that meets the adequacy requirement. This formulation tends to underestimate risk because of its limited stochastic coverage.
@@
  • SPN: Uses the fitted SPN (using the available data sampled for the specific iteration) to directly estimate the probability of adequacy violations. This estimation cannot be the input of an optimizer, but it is still a point of reference for what the best solution the SPN-based strategies could yield.
  • SPN-Max: This value follows the estimation proposed in [nemecek2024generating] to have an MILP-compatible function. This approach enables the embedding of SPNs within a tractable optimization pipeline, allowing for the explicit enforcement of chance constraints during decision-making.
  • SPN-Piecewise: This is an improved formulation of SPN using MILP constraints, using piece-wise linear functions to approximate the exponential and logarithm functions. For each sum node $n^{\Sigma}$, the maximal input value is subtracted from all values for numerical stability; then, we approximate the exponentials on the range $\ln 0.001$, 0] where values are numerically relevant (values below 0.001 are rounded to 0). Then, a logarithm approximation of the sum of exponential outputs is taken, where the logarithm is approximated on [1, $|\mathrm{ch}(n^{\Sigma})|$] to cover all possible values. Each function is approximated using five linear segments. The output of the logarithm is finally summed with the subtracted maximal value, providing a tighter approximation of the true value.
@@
This experiment yields important insights that validate our method, while also highlighting key considerations for future research. First, it illustrates the progressive improvement of the SPN’s reliability with larger dataset sizes. A limited number of samples allows fitting an SPN roughly; it will learn some high-level patterns that might allow finding a feasible region to make a decision, but as it receives more data, it will undergo a process of refinement that could damage the quality of the estimation (like in the case of $50\%$). The SPN-based approximations depend directly on how well the SPN is fitted, which will be posteriorly encoded in the MILP. However, our test showed that the error introduced by the approximation of the sum node might have a regularization effect (causing the decision to be conservative), which, as a consequence, keeps the decision consistently within the allowance constraint throughout the different data availability scenarios, which, in our experiment, is essential.
@@
Comparing the SPN approach with GEP10 and GEP50 reveals several advantages to using SPNs. In our experimental setup, traditional GEP approaches heavily depend on the quality of the selected scenarios, which we test by selecting them from the available samples at each iteration of our experiment. When choosing a few scenarios, there is more uncertainty in fulfilling the adequacy constraint, as evidenced in Figure 4, where GEP10 shows large variability among the iterations and provides no guarantees of preventing shortfall. Switching to GEP50, considering a larger set of reference scenarios improves the consistency of the decision; however, it does not offer certainty about respecting the adequacy either. In both cases, the quality of the available scenarios becomes a matter.
@@
We noticed that GEP10 exhibits high variability in the cost of the selected solution, which is a consequence of its dependence on scenario selection and the limitation on the number of scenarios that can be considered ($K$). The case of GEP50 shows an expected outcome: solutions would tend to be more expensive than the optimal one, but they show the lowest deviation from the optimal cost. However, the average solution lies on the boundary of adequacy, meaning that any error in the selection or simulation could cause the solution to move outside the safe area. Finally, the SPN methods tend to be conservative in their decisions, a result of the quality of the fitting process and the error introduced by the approximation. SPN methods demonstrate greater consistency across various data availability scenarios and random samples. This is a clear advantage, considering that the operation cost component $C^{\mathrm{op}}(x,\xi)$ in (4) would be negatively affected by shortfall events. Thus, we argue that TPM-based methods for solving the chance-constrained Capacity/Generation/Storage expansion problem offer a safer and more cost-effective (in the long term) alternative to traditional methods.
@@ 4.3 Training Details
We trained our SPNs using the SPFlow library [Molina2019SPFlow], which supports automatic structure learning and exact inference for tractable probabilistic models through queries. The structure learning procedure relies on hierarchical clustering and recursive variable partitioning to identify conditional independencies within the data. We optimized the hyperparameters using WandB [wandb] through Bayesian search over 500 trials, per percentage of the total dataset, with model selection based on ground-truth log-likelihood on the validation set. We standardized and discretized features depending on the leaf node type. Individual runs required from a few seconds to several minutes, depending on the dataset size and model complexity. Early stopping based on validation likelihood was employed to prevent overfitting. To ensure full reproducibility, we plan to make all code, configuration files, and instructions publicly available, thereby facilitating the replication of the experiments.
@@ 5 Conclusion
This work presents a probabilistic framework for investment planning in power systems under uncertainty, based on Sum–Product Networks (SPNs), by embedding these SPNs within a mixed-integer linear programming (MILP) formulation. Based on existing work on Tractable Probabilistic Models (TPMs), we demonstrated that probabilistic reasoning can be incorporated into chance-constrained optimization problems without explicit scenario enumeration, thereby redefining the traditional capacity/generation/storage scenario expansion problem by integrating a probabilistic adequacy constraint.
Our experimental results on a representative generation–storage planning case study indicated that SPN-based models can accurately capture adequacy probabilities and maintain reliable performance even under limited data availability. Compared to traditional scenario-based methods, the SPN formulation yielded more consistent and conservative investment decisions, mitigating shortfall risk at moderate computational cost. These findings highlight the potential of TPMs as a foundation for scalable, data-driven, and reliability-aware expansion planning in future low-carbon power systems that rely on stochastic, non-dispatchable sources. This work shows that TPMs offer a promising and interpretable approach to integrating uncertainty quantification and optimization in long-term power system planning.
@@ 6 Limitations
While the proposed framework exhibits strong potential, several significant limitations are worth highlighting, which could facilitate further research. First, the SPN linearization introduces an approximation that may lead to conservative or locally imprecise density estimation, yielding a calibration problem and potentially introducing overconfidence when multiple probabilistic paths contribute significantly to the overall likelihood. Future work will explore alternative formulations, other probabilistic circuits, or hybrid approaches to reduce this approximation error. Additionally, the quality of probabilistic inference depends on the representativeness of the training data. Real-world datasets may exhibit bias or insufficient coverage of rare events, which could affect reliability estimation. Incorporating active learning or uncertainty-aware sampling could improve robustness in such settings. Although the TPM formulation scales more favorably than scenario-based methods, embedding large SPNs within MILP solvers can pose computational challenges for full-scale systems if the learned structure becomes increasingly complex.